50 mL of yellowish liquid collected.
Monthly Archives: August 2014
140829 UCSC
full bottle collected. Split 15 mL for anion analysis, 100 mL were filtered with a bottle top 0.2 micron filter into a glass bottle then transferred into an I-Chem jar. This was then acidified, as was the unfiltered 125 mL in the original jar.
140828 ISB(UCSC)
Robert arrived at ISB at 9:15. The doors were open but no sample had been collected. The top plate, drip tray and the strand frames were removed and washed with 409 and DI. The inside of the collector was also rinsed with DI water. The system was reset for another sample.
140827 UCSC
Kona got to the roof of ISB at 9 am. There was <5 mL of water in the collecting jar which was discarded. The apparatus was cleaned in place with 409 and DI, then the system was reset for fog collection.
140826
Robert arrived at LML at 9:15. The doors were open but no sample had been collected. The top plate, drip tray and the strand frames were removed and washed with 409 and DI. The inside of the collector was also rinsed with DI water. The system was reset for another sample.
140824 LML
Robert arrived at LML at 9:15. The front door to the collector had not opened but the back one had. The solenoid was adjusted in an attempt to remedy the problem and then tested successfully several times.
140825 UCSC
No fog sample overnight. Heavy high stratus. Not burned off by noon, no signs of letting up today.
140825 LML
No fog sample overnight. Front door did not open. Adjusted it (increased the tension). Latches are working well. Should be good to go.
140823 UCSC
High stratus over night with no wetness on the ground in the morning. Arrived at 11 am to ISB greenhouse and there was about 20 mL of clear sample in the bottle. Also the funnel/rain collector had about 3 mL in in it. I kept the sample from fog collector and discarded the “rain” sample. Cleaned the collector by rinsing in place. Attached new bottle. Reset program.
ISB Observations 8/22
Jeremiah went to the ISB roof this morning. A 75mL sample (S-UCSC-140822) was present, and a blank (B-UCSC-140822) was taken after cleaning the collector. The collector is in very good condition. The sample collected this morning contained a thin layer of what looked like dirt or soil at the bottom.